OpenStudy (anonymous):

Simplify the product. (x+4)(x+3) A.x^2-7x-12 B.x^2+x-12 C.x^2-x-12 D.x^2-x-12-1

OpenStudy (hoblos):

(x+4)(x+3)= x^2 +3x+4x+12 = x^2 +7x+12

OpenStudy (hoblos):

its not one of the options though!! check them

OpenStudy (anonymous):

Wait , no ! My mistake the question was (x-4)(x+3)

OpenStudy (hoblos):

(x-4)(x+3) = x^2 -x -12